OpenStudy (anonymous):

Find the exact value of cos(θ/2) given cosθ=1/6; 0 degrees<θ<90 degrees PLEASE HELP

OpenStudy (anonymous):

cos(theta)=2*cos^2(theta/2)-1 = sqrt(0.5*(cos(thta)+1) =sqrt(7/12)

OpenStudy (anonymous):

:( That's what I got too but its not one of the options....

OpenStudy (anonymous):

My options are: a. sqrt(35)/18 b. -17/18 c. sqrt(15)/6 d. sqrt(21)/6